### v2

# How (Mis)calibrated is Your Federated CLIP and What To Do About It?

Abstract
While vision-language models like CLIP have been extensively studied, their calibration, crucial for reliable predictions, has received limited attention. Although a few prior works have examined CLIP calibration in offline settings, the impact of fine-tuning CLIP in a federated learning (FL) setup remains unexplored. In this work, we investigate how FL affects CLIP calibration and propose strategies to improve reliability in this distributed setting. We first analyze Textual Prompt Tuning approaches and show that they degrade calibration metrics when operating under FL. We also evaluate existing in-training calibration techniques across four global aggregation methods, finding that they provide limited improvements. Our results suggest that the key challenge lies not only in how we aggregate or calibrate, but in which components we choose to fine-tune. Motivated by this insight, we propose $\text{FL}^{2}\text{oRA}$, a straightforward LoRA-based approach that naturally improves calibration in FL, and we analyze the factors behind its effectiveness. Experiments on multiple benchmarks demonstrate that $\text{FL}^{2}\text{oRA}$ consistently produces well-calibrated models, reducing the need for explicit calibration procedures. Codes are available at https://github.com/mainaksingha01/FL2oRA.
1 Introduction
Reliability is a critical consideration when deploying deep learning (DL) models in production scenarios. We aim to develop DL models that are not only highly accurate but also reliable. Reliability has many facets, including robustness to covariate shift [csurka2017domain], model calibration [guo2017calibration], and generalization to out-of-distribution data [yang2024generalized]. Of particular interest, model calibration is crucial because it ensures that a model’s predicted probabilities accurately reflect the true likelihood of correctness, enabling reliable decision-making [pavlovic2025understanding]. As powerful DL models (e.g., CLIP [radford2019language]) are increasingly being deployed in high-stake applications, such as healthcare and autonomous systems, ensuring proper calibration is crucial.
Model calibration is a widely studied topic in machine learning [wang2023calibration, geng2024survey]. Prior works have shown that although vision-language models, such as CLIP, are reasonably well-calibrated for zero-shot image classification [minderer2021revisiting], fine-tuning them on downstream tasks often leads to miscalibration [dor, wang2024open]. This is concerning because, due to its high generalizability, CLIP is often fine-tuned on small target datasets, which increases the risk of miscalibration. To alleviate miscalibration in downstream tasks, either post-hoc calibration techniques (e.g., temperature scaling [guo2017calibration]) or in-training calibration techniques (e.g., DCA [dca], DOR [dor]) have been used. However, fine-tuning is not only limited to the offline setting, where a dataset is available in a centralized location [coop, kgcoop, maple], but also in federated learning (FL) [bai2024diprompt, promptfl, qiu2024federated, sun2024towards, fedotp, fedtpg, fedpgp, fedpha], which is a more realistic setup. To our surprise, despite the practical relevance and progress in FL, no prior work has investigated how fine-tuning CLIP affects model calibration in the FL setting.
Figure 1: Comparison of reliability diagrams and calibration errors. Expected Calibration Error (ECE $\downarrow$) [ece] is reported on the OxfordPets [oxfordpets] dataset on a centralized training or offline (left) and a non-IID personalized federal learning (FL) setting (right).
Table 1: Overview of works focusing on model calibration. While previous works have studied calibration of CLIP, no work has considered how fine-tuning CLIP in the FL setting affects calibration. Our work explores this direction.
Method CLIP Calibration Fine-tune FL
Minderer et al. [minderer2021revisiting]
Wang et al. [dor]
Chu et al. [nucfl]
Our work
In this work, we make an attempt to answer the question: how much calibrated is CLIP when fine-tuned in a Federated Learning setup? Towards that goal, we first analyze if the trend of miscalibration, observed in the offline CLIP fine-tuning setting [dor, wang2024open], also holds true in the FL setting. As shown in Fig. 1, CoOp [coop], a textual-prompt tuning approach, when fine-tuned in the FL setup (PromptFL [promptfl]) is even more miscalibrated. Imposing regularization through in-training calibration (e.g., with DCA [dca]) to both the offline and FL setting alleviates miscalibration by some extent. This observation conforms with the findings of [nucfl], where a similar behaviour is shown in the FL setup.
However, as shown in Tab. 1, the work in [nucfl] only experimented with the classical ResNet family of models trained from scratch and did not explore a multimodal setting. Differently, in our study we are specifically interested in analyzing the miscalibration of fine-tuned CLIP models in the FL setup, which exhibits substantially different dynamics from standard Convnets [minderer2021revisiting]. In Sec. 4, we dive deep into this issue, by investigating existing methods for improving calibration, and our empirical analyses show that improving calibration of CLIP-based methods in the FL setup is more challenging than the offline setting.
Core findings of our work Fine-tuning CLIP in the FL setting is more prone to miscalibration than in the centralized setting Fine-tuning strategy affect a lot the calibration, with LoRA being more effective than other PEFT methods Explicitly optimizing for calibration rarely produces systematic improvements, and often can be detrimental
As an alternative, we strive to find a simple approach that not only preserves the accuracy of the model, but also ensures a well-calibrated model without introducing additional regularization. To this end, we deviate from the common approach of textual prompt-tuning in FL [promptfl, diprompt, fedotp], and adopt low rank adaptation (LoRA [lora]) as a means to fine-tune CLIP. We call our approach $\text{FL}^{2}\text{oRA}$ (Federated Learning with LoRA). As shown in Fig. 1 (right), $\text{FL}^{2}\text{oRA}$ alone is better calibrated than PromptFL + DCA, without the need for any explicit regularization on the network. With extensive experiments on three FL settings that encompass generalization to new classes and domains and their combination, we show that fine-tuning CLIP with LoRA alone is highly effective in preserving model calibration in the FL setting, and adding extra regularization to improve calibration brings only a marginal improvement. Note that the goal of this work is not to introduce a novel architecture for FL, but to fill the gap in the literature by empirically analyzing calibration in federated CLIP fine-tuning and proposing a simple, while effective approach to preserve calibration.
Contributions. In summary: (i) We are the first to empirically study the impact of fine-tuning CLIP on the model’s calibration in the FL setup. (ii) We present $\text{FL}^{2}\text{oRA}$, a simple and effective approach, that departs from the dominant paradigm of fine-tuning CLIP with textual prompt-tuning. (iii) We conduct extensive experiments across diverse FL scenarios and show that our LoRA-based federated fine-tuning consistently provides a strong baseline, potentially removing the need for additional regularization for calibration.
2 Related Work
Federated learning enables collaborative model training across decentralized clients without sharing local data. A core challenge is client heterogeneity, e.g., non-IID data distributions, which can significantly degrade performance. Prior work primarily focuses on designing robust aggregation and optimization strategies to address this issue. Representative methods include FedAvg [fedavg], which averages client updates, FedProx [fedprox], which introduces a proximal term to mitigate client drift and FedNova [fednova], which normalizes updates to improve convergence under heterogeneous objectives. Broadly, these approaches aim to improve the accuracy and convergence of the aggregated global model. In contrast, we investigate a relatively underexplored aspect of FL: its impact on model calibration in Vision Language Models (VLMs).
Federated fine-tuning of CLIP. When applied to large VLMs such as CLIP [radford2019language], FL introduces substantial computational and communication overhead due to the need to fine-tune millions of parameters. To mitigate this, prompt tuning [coop, kgcoop, vpt, maple] has emerged as an efficient alternative for adapting CLIP. Recent works extend prompt tuning to the federated setting [bai2024diprompt, promptfl, qiu2024federated, sun2024towards, fedotp, fedtpg, fedpgp, fedpha]. For example, PromptFL [promptfl] regularizes client-specific prompts, while FedOTP [fedotp] and FedTPG [fedtpg] improve communication efficiency and personalization via orthogonal or task-conditioned prompt generation. FedCLIP [fedclip] decouples global and local adaptation of the text and image encoders, and methods such as FedTPG [fedtpg] and FedMVP [fedmvp] dynamically generate prompts using auxiliary information (e.g., class names or visual attributes).
In contrast, we adopt a Parameter-Efficient Fine-Tuning (PEFT) approach, i.e., LoRA [biderman2024lora], to adapt CLIP in the federated setting. While FL with LoRA has been studied for LLMs [yang2025federated], prior work focuses mainly on accuracy. To our knowledge, we are the first to study calibration in CLIP-based federated image classification.
Model calibration [pavlovic2025understanding] measures how well predicted confidence scores reflect true correctness likelihood. Calibration methods are generally divided into post-hoc and training-time approaches. Post-hoc methods, such as temperature scaling [guo2017calibration] and Dirichlet calibration [kull2017beta], adjust output probabilities after training without modifying model parameters, offering simplicity but limited flexibility. In contrast, training-time methods, including label smoothing [muller2019does], focal loss [lin2017focal], DCA [dca], and MDCA [hebbalaguppe2022stitch], modify the objective to encourage well-calibrated predictions, often at the cost of additional hyperparameter tuning.
For VLMs, zero-shot CLIP is known to be well-calibrated [minderer2021revisiting], but fine-tuning can degrade calibration, motivating approaches such as DOR [dor], DAC [wang2024open], and CalShift [khan2025confidence]. Most prior studies focus on offline settings. In FL, only a few works [peng2024fedcal, nucfl] address calibration, primarily with fully fine-tuned CNNs. To our knowledge, calibration in CLIP-based federated learning remains unexplored.
3 Preliminaries
In this section, we discuss the preliminaries related to federated learning, federated fine-tuning of CLIP and model calibration.
Federated learning. For the task of image classification, the FL scheme consists of training data distributed among $N$ clients and each client $n$ possess a dataset $\mathcal{D}_{n}=\{\mathbf{x}^{(i)},y^{(i)}\}^{\mid\mathcal{D}_{n}\mid}_{i=1}$, where $\mathbf{x}^{(i)}$ denotes the $i$-th input image and $y^{(i)}$ denotes the class label. Ideally, we want to train a model by minimizing the empirical risk on the union of datasets of all the clients, $\mathcal{D}=\cup_{n\in N}\mathcal{D}_{n}$. However, due to the distributed nature of the datasets in the FL setting, we can only train a model $f_{\theta}$ on the local data of the client, i.e., conduct local training, as:
$\min_{\theta}\mathcal{L}(\theta)=\frac{1}{|\mathcal{D}_{n}|}\sum^{|\mathcal{D}_{n}|}_{i=1}\mathcal{L}_{\text{ce}}(f_{\theta}(\mathbf{x}^{(i)}),y^{(i)}),$ (1)
where $\theta$ are the weights of a client model and $\mathcal{L}_{\text{ce}}$ is cross-entropy loss.
To aggregate knowledge (or global aggregation), the FL setting involves clients communicating the weights of the trained model $f_{\theta}$ to the server through multiple rounds of communication. In each round $t$, the server aggregates client weights via a weighted average, or, more concretely, using FedAvg [fedavg] as:
$\theta_{t}=\sum^{N}_{n=1}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\theta^{(n)}_{t}$, (2)
where $\theta^{(n)}_{t}$ is the weights of the $n$-th client. After aggregation, the server sends back $\theta_{t}$ to the clients, which is used as a starting point for further local training.
In this work, we operate in a non-IID FL setup [hsu2019measuring], where each client $n$ can have varying number of classes, and the classes overlap among clients. This is more challenging than the prior CLIP-based FL literature [fedtpg] that assumes the classes to be disjoint among the clients.
Federated fine-tuning of CLIP. Textual prompt tuning (TPT), i.e., CoOp [coop], is a commonly used fine-tuning approach to improve the classification accuracy of CLIP on downstream tasks. For fine-tuning, TPT learns a set of prompts while keeping the weights of the pretrained CLIP image encoder $\mathcal{E}_{v}$ and text encoder $\mathcal{E}_{t}$ frozen. In detail, TPT replaces handcrafted prompts (e.g., “a photo of a [CLS]”) at the input of the text encoder $\mathcal{E}_{t}$ with a set of lightweight learnable vectors (or tokens) $\mathcal{T}=\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{M}\}$, where $M$ denotes the length of the tokens. Text features are computed with the prompt $\mathbf{t}=\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{M},[\text{CLS}]\}$, and the image features are computed from the image $\mathbf{x}$, using the respective encoders. The prompts $\mathcal{T}$ are learned using cross-entropy loss, as in Eq. (1).
Since the prompts $\mathcal{T}$ in TPT are lightweight and incur low communication overhead, it has become a dominant approach for fine-tuning CLIP in the FL setting [promptfl, diprompt, fedclip]. Each client trains prompts $\mathcal{T}$ on the local client data and communicates them with the server. The server aggregates the prompts $\mathcal{T}$ from different clients using Eq. (2), and sends back the aggregated prompts to the clients. TPT-based FL methods differ mostly in how the prompts are designed, trained locally, and aggregated on the server.
While TPT is not the only method for fine-tuning CLIP in FL, it remains the most widely used. Related work has explored Adapters for FL of CLIP [fedclip] and LoRA for fine-tuning LLMs in multi-task and federated settings [bohdal2025efficient, yang2025federated].
Model calibration. In addition to higher accuracy, we also want a model to be well-calibrated. That is, we want the predicted class probabilities of the model to be an accurate representation of the model’s accuracy [guo2017calibration]. Formally, a model is said to be well-calibrated (or confidence-calibrated), if for a confidence of $c$, the model is correct $c$ proportions of the time:
$\mathbb{P}(y=\arg\max(\hat{p}(\mathbf{x}))\mid\max(\hat{p}(\mathbf{x}))=c)=c\quad\forall c\in[0,1],$
[…]
$\text{ECE}=\sum^{G}_{g=1}\frac{\mid b_{g}\mid}{|\mathcal{D}|}\mid\text{acc}(b_{g})-\text{conf}(b_{g})\mid,$
where $\text{acc}(b_{g})$ and $\text{conf}(g_{g})$ denotes the average accuracy and average confidence of the model for the samples falling in the bin $b_{g}$, respectively. A well-calibrated model will have a low ECE.
In centralized (offline) settings, prior work showed that while zero-shot CLIP is well-calibrated in few-shot classification [minderer2021revisiting], fine-tuning can degrade calibration [wang2024open, dor]. To mitigate this, methods introduce regularization (e.g., constraining text encoder divergence [dor]) during fine-tuning. However, the impact of fine-tuning CLIP on calibration in FL remains unexplored.
4 Empirical Evaluation
4.1 Experimental setup
Benchmarks and settings. For extensive evaluation, we use a collection of 10 benchmarks: CIFAR-10, CIFAR-100 [cifar], OfficeHome [officehome], PACS [pacs], VLCS [vlcs], Food101 [food], DTD [dtd], Caltech101 [caltech], Flowers102 [oxfordflowers], and OxfordPets [oxfordpets]. We follow the non-IID FL setting, where the clients have training data of variable number of classes and the classes can be shared among the clients, i.e. overlapping classes. Following previous work [cao2023knowledge, shamsian2021personalized], we simulate data heterogeneity by partitioning the data among the clients by setting the parameter $\beta=0.5$ of a symmetric Dirichlet distribution.
We organize the benchmarks to simulate heterogeneous FL scenarios: (i) an in-distribution setting where a single dataset (e.g., CIFAR-100) is split across clients, (ii) a domain generalization setting where different domains (e.g., OfficeHome) are assigned to different clients, and (iii) a base-to-new generalization setting where some classes are held out during training. Additional details are provided in the Appendix.
Baselines. We evaluate several TPT methods for CLIP-based image classification: (i) PromptFL [promptfl], which extends CoOp [coop] to FL; (ii) KgCoOp [kgcoop], which adds a regularization term to CoOp, along with its federated variant, FedKgCoOp; (iii) state-of-the-art TPT methods for personalized FL: FedPGP [fedpgp], FedOTP [fedotp], and FedPHA [fedpha]. For off-the-shelf calibration, we use temperature scaling [guo2017calibration] (post-hoc method) and two training-based methods: DCA [dca] and MDCA [mdca].
Implementation Details. We use a frozen CLIP ViT-B/16 backbone. We set the Dirichlet parameter to $\beta=0.5$, run 50 communication rounds with 1 local epoch, and use 100% participation rate (10% for CIFAR-10/100). For base-to-new generalization, we use 16 shots per class across 10 clients; for domain generalization, two clients per domain; and for in-distribution settings, 100 clients per dataset. We follow the standard optimizer setup with SGD [sgd], a learning rate of $1\times 10^{-3}$, warm-up $1\times 10^{-5}$, and batch size 32.
Evaluation metrics. To compare methods, we compute: classification accuracy (Acc $\uparrow$), Expected Calibration Error (ECE $\downarrow$) [guo2017calibration], Maximum Calibration Error (MCE $\downarrow$) [guo2017calibration], and Adaptive Calibration Error (ACE $\downarrow$) [nixon2019measuring]. We report performances in the personalized FL setting [fedotp, fedpgp, fedpha], where the metrics are computed at each client and then averaged across clients.
Table 2: Comparison of offline CLIP fine-tuning with federated CLIP fine-tuning. Results have been reported on the in-distribution setting using datasets: CIFAR-10 and CIFAR-100. Best number is in bold.
Method Offline setting (w/o calib) Offline setting (w/ calib: DCA [dca]) FL setting (w/o calib)
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
CIFAR-10
CoOp [coop] 93.99 1.00 0.56 0.89 93.89 1.20 0.62 0.87 92.42 4.94 1.64 3.94
KgCoOp [kgcoop] 92.17 2.70 0.61 2.70 92.37 2.21 0.46 2.21 90.48 6.53 1.88 5.79
VPT [vpt] 96.93 1.23 0.64 0.90 96.98 1.45 0.75 0.94 95.29 3.53 1.22 2.52
CIFAR-100
CoOp [coop] 77.86 1.93 0.57 1.77 77.85 1.83 0.73 1.71 72.43 9.27 2.43 8.66
KgCoOp [kgcoop] 74.73 7.83 1.13 7.83 74.82 5.30 0.80 5.30 70.68 11.20 2.95 11.05
VPT [vpt] 80.86 1.45 0.41 1.35 80.82 1.47 0.39 1.46 77.83 7.94 2.39 8.00
4.2 How does Federated Learning affect calibration?
In this section, we compare the calibration error of TPT-based methods (e.g., CoOp, KgCoOp) in both the offline and FL setting. Our goal is to understand whether the miscalibration of TPT-based methods, observed in [dor] in the offline CLIP fine-tuning setting, also holds true in the FL setting. Tab.˜2 reports the performance of TPT methods: (i) in the offline setting without calibration, (ii) in the offline setting with calibration (i.e., with DCA [dca]), and (iii) in the FL setting without calibration. In addition, we consider the visual prompt tuning method, VPT [vpt], for comprehensiveness. The FL experiments were performed in the in-distribution setting.
On average across metrics and datasets, offline prompt-tuning methods exhibit higher calibration error than when combined with a calibration technique. This aligns with prior empirical findings [dor]. A previously unexplored question is: to what extent does fine-tuning CLIP in a distributed setting worsen calibration? As shown in the last column of Tab. 2, the same methods exhibit substantially higher calibration error in FL across all metrics. While prior work [nucfl] reports similar trends, it is limited to ResNets trained from scratch. In contrast, we study CLIP fine-tuning in FL, which is particularly relevant given its widespread downstream use.
In summary, the results in Tab. 2 indicate that fine-tuning CLIP in the FL setting consistently introduces miscalibration, mirroring observations for ResNets [nucfl]. Therefore, assessing CLIP-based FL methods solely based on accuracy, as done in several recent works [fedotp, fedpha, fedpgp], provides an incomplete evaluation of their performance.
Table 3: Comparison of federated CLIP fine-tuning methods. Results have been reported on the in-distribution setting using datasets: CIFAR-10 and CIFAR-100. Best number is in bold, and second best is underlined.
Method FL setting (w/o calib) FL setting (w/ calib: DCA [dca]) FL setting (w/ calib: MDCA [mdca])
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
CIFAR-10
PromptFL [promptfl] 92.42 4.94 1.64 3.94 92.12 4.98 1.73 3.92 92.31 5.07 1.67 3.95
FedPGP [fedpgp] 91.28 5.89 1.78 5.17 91.54 5.65 1.75 4.69 90.98 5.68 1.79 5.06
FedOTP [fedotp] 94.38 6.03 1.74 5.37 94.30 5.41 1.59 4.53 94.24 5.83 1.77 5.17
FedPHA [fedpha] 93.72 4.59 1.47 3.52 93.53 4.81 1.46 3.41 93.55 4.78 1.53 3.38
$\text{FL}^{2}\text{oRA}$ (Ours) 97.04 2.99 1.12 2.06 97.00 2.99 1.15 1.87 96.92 3.03 1.15 1.96
CIFAR-100
PromptFL [promptfl] 72.43 9.27 2.43 8.66 72.07 8.92 2.34 8.77 72.02 9.18 2.45 8.85
FedPGP [fedpgp] 71.13 9.91 2.59 9.78 71.36 10.38 2.73 10.14 71.88 9.82 2.66 9.65
FedOTP [fedotp] 74.81 14.44 3.63 14.30 74.63 12.67 3.27 12.52 74.73 14.42 3.60 14.45
FedPHA [fedpha] 74.23 8.91 2.33 8.61 75.30 8.06 2.20 7.63 75.34 8.65 2.39 8.11
$\text{FL}^{2}\text{oRA}$ (Ours) 83.70 6.70 2.05 6.09 83.65 6.64 2.12 6.20 83.73 6.82 2.01 6.13
4.3 Can we improve calibration?
Given that prompt tuning methods worsen calibration in FL, a natural question is whether calibration can be meaningfully improved and how. To answer this, we systematically evaluate state-of-the-art FL baselines in three settings: in-distribution (Tab. 3), domain generalization (Tab. 4), and base-to-new generalization (Tab. 5), with and without calibration regularization.
Table 4: Comparison of federated CLIP fine-tuning methods. Average performance have been reported on the domain generalization setting using the datasets: OfficeHome, PACS and VLCS. Best number is in bold, and second best is underlined.
Method FL setting (w/o calib) FL setting (w/ calib: DCA [dca]) FL setting (w/ calib: MDCA [mdca])
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
PromptFL [promptfl] 91.25 2.88 1.21 2.79 91.32 3.07 1.22 2.87 91.32 2.91 1.21 2.77
FedPGP [fedpgp] 92.71 2.67 0.88 2.24 93.93 2.28 0.68 2.06 94.04 2.29 0.90 1.90
FedOTP [fedotp] 91.07 9.34 2.36 9.20 91.54 7.37 1.90 7.27 91.88 9.26 2.40 9.24
FedPHA [fedpha] 92.87 2.29 0.81 1.96 93.93 2.17 0.64 2.02 94.06 2.23 0.94 1.83
$\text{FL}^{2}\text{oRA}$ (Ours) 93.11 2.37 0.79 1.83 93.13 2.09 0.72 1.95 93.07 2.24 0.77 1.64
Table 5: Comparison of federated CLIP fine-tuning methods. Average performance have been reported on the base-to-new generalization setting using the datasets: Food101, DTD, Caltech101, Flowers102, and OxfordPets. Best number is in bold, and second best is underlined.
Method FL setting (w/o calib) FL setting (w/ calib: DCA [dca]) FL setting (w/ calib: MDCA [mdca])
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
PromptFL [promptfl] 82.98 4.11 1.31 3.82 84.25 3.97 1.30 3.62 84.62 4.06 1.23 3.73
FedPGP [fedpgp] 81.09 6.17 2.07 5.97 80.90 5.82 1.82 5.69 81.51 4.93 1.65 4.80
FedOTP [fedotp] 29.52 21.88 5.03 21.90 28.63 26.57 7.34 26.63 28.53 25.35 6.25 25.32
FedPHA [fedpha] 54.97 16.03 4.64 15.69 54.56 17.93 5.53 17.38 52.34 20.55 6.63 19.86
$\text{FL}^{2}\text{oRA}$ (Ours) 86.25 3.91 1.41 3.20 86.08 4.21 1.61 3.32 86.34 4.10 1.55 3.35
Effect of calibration techniques. A seemingly straightforward way to improve calibration is to impose calibration regularization (e.g., with DCA [dca] or MDCA [mdca]) during federated fine-tuning. We applied these techniques and report results of our experiments in Tab.˜3. Applying DCA or MDCA to FL prompt-tuning methods yields only modest improvements in ECE and ACE for certain configurations (e.g., PromptFL and FedPHA on CIFAR-100), while many methods remain substantially miscalibrated. For example, on CIFAR-100, FedOTP still exhibits ECE above 12% with DCA and above 14% with MDCA. A similar pattern appears on domain generalization and base-to-new generalization (Tabs. 4 and 5). DCA and MDCA slightly improve calibration on FedPGP and FedPHA in some cases, yet other methods (again notably FedOTP) stay severely miscalibrated, with ECE often exceeding 7–9%. Overall, we observe that in-training calibration in TPT for FL rarely produces large, systematic improvements and sometimes even makes it more miscalibrated.
Effect of global aggregation. Next, we test whether stronger FL optimizers can alleviate miscalibration. As shown in Tab.˜6, replacing FedAvg with FedProx, FedDyn or FedNova produces no consistent improvement in either accuracy or calibration. The calibration metrics of these optimizers closely track those of FedAvg and typically differ only within run-to-run variability; in several cases, FedAvg is slightly better. We conclude that calibration issues of TPT for FL are not primarily driven by the choice of global aggregation algorithm.
Improving calibration through LoRA fine-tuning. Most existing FL approaches for CLIP fine-tuning rely on textual prompt tuning. To provide a different perspective, we ask: could the miscalibration issue stem from what we choose to fine-tune? To find an answer, we propose to replace text prompts with LoRA for fine-tuning CLIP. LoRA [lora] augments linear layers of CLIP’s vision and/or text encoders with low-rank adapters. LoRA applies to any linear transformation with weights $W\in\mathbb{R}^{m\times n}$ by expressing the update $\delta W=\alpha AB$ in low rank matrices $A\in\mathbb{R}^{m\times r},B\in\mathbb{R}^{r\times n}$, where $r$ is the LoRA rank and $\alpha$ is a fixed scaling factor (typically $\alpha=1/r$). We apply LoRA to the Q, K, and V projection layers of all blocks in both encoders during fine-tuning, and call our method $\text{FL}^{2}\text{oRA}$. The implementation details of $\text{FL}^{2}\text{oRA}$ are in the Appendix.
Table 6: Comparison of different global aggregation algorithms. Results have been reported on the in-distribution setting without applying any calibration using CIFAR-100 dataset. Best number is in bold, and second best is underlined.
Method FedAvg[fedavg] FedProx[fedprox] FedDyn[feddyn] FedNova[fednova]
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
PromptFL 72.43 9.27 2.43 8.66 73.12 8.16 2.24 7.29 73.28 8.34 2.35 7.56 72.89 9.54 3.27 8.89
FedKgCoOp 70.68 11.20 2.95 11.05 71.56 11.08 2.86 8.57 71.78 11.45 2.39 7.34 70.38 12.38 3.21 8.32
FedMaPLe 81.73 7.53 2.20 6.93 82.68 7.84 2.34 7.05 82.07 7.08 2.19 7.06 81.38 8.26 2.93 8.19
$\text{FL}^{2}\text{oRA}$ (Ours) 83.70 6.70 2.05 6.09 83.55 6.86 2.14 6.15 83.42 6.79 2.15 6.15 82.89 7.03 2.56 6.67
From Tab. 3, we observe that $\text{FL}^{2}\text{oRA}$ clearly outperforms all the prompt-tuning baselines in all three ECE/MCE/ACE metrics, even without needing additional regularization. For example, on CIFAR-100 without calibration, $\text{FL}^{2}\text{oRA}$ improves accuracy from 74.81% (FedOTP) to 83.70% and reduces ECE from 14.44% (FedOTP) and 8.91% (FedPHA) down to 6.70%. On the more challenging base-to-new setting (Tab. 5), where all methods exhibit worse calibration and additional calibration has some effect, $\text{FL}^{2}\text{oRA}$ demonstrates all-round performance across all configurations. On the domain generalization benchmark (Tab. 4), the gap between prompt-tuning methods and $\text{FL}^{2}\text{oRA}$ becomes smaller, as DG poses a less severe adaptation challenge for CLIP: the input resolution and the label set remain unchanged, and only the image distribution shifts. Nevertheless, $\text{FL}^{2}\text{oRA}$ still achieves the highest accuracy without calibration (93.11%) and, after applying DCA or MDCA, provides the best calibration trade-off, attaining the lowest ACE with DCA (1.95) and MDCA (1.64).
Taken together, these results indicate that CLIP with parameter-efficient LoRA fine-tuning or $\text{FL}^{2}\text{oRA}$ naturally produces well-calibrated models in FL, substantially reducing the need for explicit calibration procedures. In contrast, CLIP-based TPT methods consistently degrade calibration in FL, and even strong in-training calibration techniques and advanced FL optimizers fail to correct this behavior fully.
4.4 What makes $\text{FL}^{2}\text{oRA}$ effective?
In this section, we provide a high-level, largely qualitative discussion of why $\text{FL}^{2}\text{oRA}$ may be particularly effective for federated CLIP fine-tuning. In the FL strategies, the final weight is a weighted average of the different learning rounds as written in Eq.˜2. Based on [izmailov2018averaging], at communication round $t$, we can decompose the output of the local model of client $n$ via a first-order Taylor expansion around the global model $\theta_{t}$:
$f_{\theta^{(n)}_{t}}(\mathbf{x}^{(i)})=f_{\theta_{t}}(\mathbf{x}^{(i)})+\bigl\langle\nabla f_{\theta_{t}}(\mathbf{x}^{(i)}),\Delta_{n}\bigr\rangle+O\bigl(\|\Delta_{n}\|^{2}\bigr),$ (3)
where $\Delta_{n}=\theta^{(n)}_{t}-\theta_{t}$ is the local deviation from the global model, and $\nabla f_{\theta_{t}}(\mathbf{x}^{(i)})$ is the Jacobian of $f$ with respect to the parameters $\theta$, evaluated at $\theta_{t}$ and input $\mathbf{x}^{(i)}$.
By multiplying Eq.˜3 by the client weight $\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}$ and summing over all clients $n$, we obtain:
$\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\,f_{\theta^{(n)}_{t}}(\mathbf{x}^{(i)})=\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\Bigl[f_{\theta_{t}}(\mathbf{x}^{(i)})+\bigl\langle\nabla f_{\theta_{t}}(\mathbf{x}^{(i)}),\Delta_{n}\bigr\rangle+O\bigl(\|\Delta_{n}\|^{2}\bigr)\Bigr].$ (4)
The left-hand side is precisely the weighted ensemble of the local model predictions, equivalent to Deep Ensemble [lakshminarayanan2017simple, yong2024spurious]. On the right-hand side, the first term simplifies as follows:
$\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\,f_{\theta_{t}}(\mathbf{x}^{(i)})=f_{\theta_{t}}(\mathbf{x}^{(i)}),$
since $\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}=1$ by definition. The linear term vanishes because:
$\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\bigl\langle\nabla f_{\theta_{t}}(\mathbf{x}^{(i)}),\Delta_{n}\bigr\rangle=\biggl\langle\nabla f_{\theta_{t}}(\mathbf{x}^{(i)}),\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\Delta_{n}\biggr\rangle=0,$
where we used the fact that $\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\Delta_{n}=\mathbf{0}$. The remainder term is therefore
$O\biggl(\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\|\Delta_{n}\|^{2}\biggr),$ which is second-order small when the local deviations $\Delta_{n}$ remain small.
Hence, the Federated Learning global model prediction $f_{\theta_{t}}(\mathbf{x}^{(i)})$ is approximately equal to the weighted Deep Ensemble: $\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\,f_{\theta^{(n)}_{t}}(\mathbf{x}^{(i)}).$
Recent work has shown that Deep Ensembles yield superior calibration compared to single models [lakshminarayanan2017simple, ovadia2019can]. In addition, in our setting, we replace full fine-tuning by LoRA [hu2022lora], so the global model corresponds to a LoRA Ensemble [wang2023lora, muhlematter2024lora]. Finally, some works have shown that LoRA Ensemble provides a computationally efficient approximation to Deep Ensemble [wenbatchensemble, muhlematter2024lora]. Based on these results, we can therefore interpret $\text{FL}^{2}\text{oRA}$ as an effective approximation to a Deep Ensemble, inheriting many of its generalization and uncertainty quantification benefits while remaining finetuning efficient.
To understand what makes $\text{FL}^{2}\text{oRA}$ different from TPT-based FL methods, where there is also an ensembling operation due to FedAvg, we take a closer at CLIP classification mechanism. In CLIP inference, the visual features are obtained from the image encoder, and the “classifier” weights are constructed from the text embeddings. In TPT-based FL methods, the ensembling operation is effectively performed only on the so-called classifier weights, and therefore can be considered equivalent to a last-layer ensemble [legate2023guiding], where only the last layer is fine-tuned and ensembled. However, without explicit mechanism to encourage diversity, such as repulsion losses, last-layer ensemble often exhibit limited functional diversity [steger2024function] compared to Deep Ensembles. In contrast, techniques such as LoRA Ensemble [muhlematter2024lora, wang2023lora], Batch Ensembles [wenbatchensemble] or Packed-Ensembles [laurent2023packed], which often operate on the level of the feature extraction backbone, have a greater member diversity.
4.5 Ablation Studies
In this section, we perform a comprehensive ablation study on the in-distribution setting using the CIFAR-10/100 dataset.
Figure 2: Ablation on the LoRA rank and placement. (left) Accuracy and (right) ECE on CIFAR-100 for different LoRA ranks $r$ and encoder placements. ‘T’ and ‘V’ denote applying LoRA to the text and vision encoders, respectively. $\text{FL}^{2}\text{oRA}$ remains stable across different ranks and consistently outperforms the best TPT baseline, even at very low ranks.
$\text{FL}^{2}\text{oRA}$ learning capacity. One may attribute the superior performance of $\text{FL}^{2}\text{oRA}$ to the larger number of trainable parameters during fine-tuning. To examine this, we analyze the effect of the LoRA rank $r$, which controls the adapter capacity to modify the base CLIP model. We also ablate the location where LoRA is applied (text and/or image encoders), which changes the total number of trainable parameters. A higher rank enables more expressive updates but increases computation and communication overhead. As shown in Fig. 2, $\text{FL}^{2}\text{oRA}$ remains stable across different ranks and consistently outperforms the best TPT baseline; even with a rank as small as 2, $\text{FL}^{2}\text{oRA}$ still achieves better performance than TPT methods.
Table 7: Effect of PEFT variants on CIFAR-10 dataset. Best number is in bold, and second best is underlined.
Method Text-encoder Image-encoder Both encoders
#Tr. Params Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ #Tr. Params Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ #Tr. Params Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
CLIP-FT [radford2019language] 38.1M 92.31 5.16 1.66 4.34 86.2M 95.57 3.87 1.41 3.33 124.3M 95.76 3.44 1.25 2.48
AdaptFormer [adaptformer] 30.8K 92.51 5.39 1.87 3.92 46.1K 96.64 3.06 1.20 2.09 76.9K 96.80 3.00 1.16 1.93
LayerNorm [layernorm] 49.1K 90.42 5.56 1.76 4.70 75.3K 96.04 3.45 1.26 2.39 124.4K 96.03 3.49 1.30 2.21
BitFit [bitfit] 24.6K 92.96 4.67 1.53 3.66 39.9K 96.09 3.08 1.19 2.25 64.5K 96.49 3.14 1.10 2.18
DoRA [dora] 92.2K 92.65 5.08 1.64 4.06 138.2K 97.23 2.99 1.12 1.90 230.4K 97.19 3.05 1.08 1.85
AdaLoRA [adalora] 73.8K 88.37 8.61 2.33 7.92 110.7K 90.19 8.26 2.31 7.71 184.5K 89.04 8.47 2.33 8.02
VeRA [vera] 36.9K 88.50 8.54 2.34 7.91 55.3K 88.77 8.55 2.37 8.19 92.2K 88.93 8.30 2.25 7.82
$\text{FL}^{2}\text{oRA}$ (Ours) 73.7K 92.64 4.99 1.68 3.77 110.6K 97.08 2.94 1.14 1.98 184.3K 97.04 2.99 1.12 2.06
Superiority of $\text{FL}^{2}\text{oRA}$ over PEFT variants. In Table 7 we show that $\text{FL}^{2}\text{oRA}$ consistently provides a stronger balance between accuracy and calibration compared to existing PEFT approaches across different encoder tuning settings. Unlike several baselines that either improve accuracy at the expense of calibration or require more trainable parameters, $\text{FL}^{2}\text{oRA}$ maintains competitive predictive performance while producing more reliable confidence estimates. This trend is particularly evident when adapting the image encoder, where $\text{FL}^{2}\text{oRA}$ achieves the most favorable calibration behavior while remaining highly accurate. Even when both encoders are updated, the method preserves this balance without introducing significant parameter overhead. These results suggest that the proposed factorized low-rank formulation improves the effectiveness of parameter-efficient adaptation by capturing task-relevant updates while avoiding the over-confident predictions commonly observed in other PEFT strategies.
(a) Accuracy
(b) ECE
Figure 3: Effect of varying communication rounds. We report (a) Accuracy and (b) ECE on CIFAR-100 dataset.
Sensitivity to communication rounds. We vary the communication rounds between 10 and 100 to study the impact of frequency of client-server communication on performance. From Fig. 3 we observe the consistent trend that all the methods benefit from more frequent communication between the clients and the server, both in terms of accuracy and ECE scores. However, $\text{FL}^{2}\text{oRA}$ achieves the best performance, with only 10 rounds of communication, even outperforming the competitors operating under maximum number of communication rounds.
Sensitivity to data distribution. We study the impact of varying $\beta$, the concentration parameter of the Dirichlet distribution, which controls the overlap of classes among clients. A low value of $\beta$ signifies a low overlap among clients (i.e., non-IID), and a high value indicates a high overlap among clients (i.e., IID). From Fig. 4 we observe that for our proposed $\text{FL}^{2}\text{oRA}$ both accuracy and ECE remain fairly stable at varying levels of heterogeneity. Although competitor methods, such as FedPGP and FedPHA, exhibit improved ECE with respect to the simple baseline ZS-CLIP, they fluctuate with different values of $\beta$.
Figure 4: Effect of varying $\beta$ of Dirichlet distribution. We report (left) Accuracy and (right) ECE on CIFAR-100 dataset.
Effect of post-hoc calibration technique. Temperature Scaling [guo2017calibration] is a widely used post-hoc calibration technique that rescales the network output (logits) by dividing them by a learned scalar parameter $\tau$, typically estimated on a validation set. Figure 5 illustrates how $\tau$ affects model calibration (ECE) on the CIFAR-10 and CIFAR-100 datasets [guo2017calibration]. The results showcase moderate temperatures yield better calibration, while too small or too large $\tau$ harms it. ZS-CLIP is highly sensitive to $\tau$ as ECE improves markedly near $\tau=0.5$ but degrades rapidly as $\tau$ increases, indicating that mild softening corrects overconfidence whereas excessive softening induces underconfident predictions. Among federated baselines, FedOTP and FedPGP tend to worsen as $\tau$ grows, suggesting a monotonic loss of calibration with over-smoothing, whereas PromptFL and FedPHA are comparatively stable with smaller ECE variation. However, our proposed $\text{FL}^{2}\text{oRA}$ maintains both the lowest and most stable ECE over a wide range of $\tau$, pointing to robustness to the temperature choice and reduced reliance on per-dataset tuning.
(a) CIFAR-10
(b) CIFAR-100
Figure 5: Effect of scaling the temperature hyperparameter ($\tau$). We report the results of (a) CIFAR-10 and (b) CIFAR-100 datasets, with ECE metric.
5 Conclusions
In this work, we investigated the an unexplored question of how fine-tuning CLIP in a FL setting affects model calibration. Our study reveals that TPT-based methods that perform reasonably in offline settings exhibit worse calibration when operating in FL. While standard in-training calibration techniques offer limited improvements, our analysis suggests that the primary driver of miscalibration is the we fine-tune CLIP. Therefore, we present $\text{FL}^{2}\text{oRA}$, and show that it naturally yield to well-calibrated models in FL, substantially reducing the need for additional calibration procedures. By exploring the link between PEFT and FL, our work opens new directions for addressing calibration in CLIP within the federated setting.
Appendix: How (Mis)calibrated is Your Federated CLIP and What To Do About It?
Appendix 0.A Experimental setup
0.A.1 Benchmarks
In our experiments, we evaluate our proposed method across three different federated learning (FL) settings using a diverse collection of benchmark datasets, as summarized in Table 8. For the in-distribution setting, we use CIFAR-10 and CIFAR-100, which are standard image classification datasets with 10 and 100 classes, respectively. Each dataset contains 50,000 training images and 10,000 test images, providing a balanced evaluation scenario where the training and test distributions are identical.
For the base-to-new generalization and domain generalization settings, we consider more challenging datasets that assess the model’s ability to transfer knowledge to unseen classes or between multiple domains. In the base-to-new setting, we choose Caltech101, Flowers102, OxfordPets, Food101, and DTD datasets in a few-shot training setup. In the domain generalization (DG) setting, we use traditional DG benchmarks having four domains each i.e. PACS (art painting, cartoon, photo and sketch), OfficeHome (art, clipart, product and realworld), and VLCS (caltech, labelme, pascal-voc and sun), providing distribution shifts.
Table 8: Statistics of the benchmarks. The table shows the statistics of the benchmarks used for empirical evaluation in three federated learning settings.
Benchmarks Domain Classes Train Test
In-distribution setting
CIFAR-10 [cifar] - 10 50,000 10,000
CIFAR-100 [cifar] - 100 50,000 10,000
Base-to-new generalization setting
Caltech101 [caltech] - 101 4,128 2,465
Flowers102 [oxfordflowers] - 102 4,093 2,463
OxfordPets [oxfordpets] - 37 2,944 3,369
Food101 [food101] - 101 50,500 30,300
DTD [dtd] - 47 2,820 1,692
Domain generalization setting
PACS [li2017deeper] Art Painting 7 1,024 614
Cartoon 1,171 704
Photo 835 502
Sketch 1,964 1,179
OfficeHome [officehome] Art 65 1,214 728
Clipart 2,191 1,298
Product 2,226 1,324
RealWorld 2,180 1,304
VLCS [vlcs] CALTECH 5 891 424
LABELME 1,672 797
PASCAL-VOC 2,127 1,013
SUN 2,067 985
Figure 6: Federated data partitioning under different Dirichlet concentration parameter $\alpha$. In each of the subplots, each row represents a client, and each cell corresponds to a data sample whose color indicates its class label. The sub figures are represented as, (a) Sort and partition setting with two fixed classes per client, (b) to (f) Non-IID splits with Dirichlet distribution, having $\alpha\!\to\!\infty$, $\alpha=100,1,0.5$, and $\alpha\!\to\!0$. Larger $\alpha$ produces nearly IID client distributions, whereas smaller $\alpha$ concentrates samples from few classes only within each client with increasingly skewed and heterogeneous partitions.
0.A.2 Federated learning settings
We describe in detail the three personalized Federated Learning: in-distribution (ID), domain generalization (DG) and base-to-new generalization.
[-] In-distribution & Domain generalization. For both settings, we distribute the classes across clients in an overlapping non-independent and identically Distributed (non-IID) manner, as similar to [fedotp]. To simulate controllable non-IID data heterogeneity across clients, we adopt the standard Dirichlet-based partitioning strategy [fedotp, cao2023knowledge, shamsian2021personalized]. Given a dataset with $C$ classes and $N$ clients, each client’s class proportion vector, $\mathbf{q}_{n}=(q_{n1},\dots,q_{nK})$ is sampled from a symmetric Dirichlet distribution [bayesian] parameterized by concentration $\alpha$:
$\mathbf{q}_{n}\sim\mathrm{Dirichlet}(\alpha\mathbf{p}),$ (5)
where $\mathbf{p}$ is the global class prior (assumed uniform in our experiments). Once the class proportions are obtained, the actual number of samples assigned to client $n$ is drawn as,
$\mathbf{c}_{n}\sim\mathrm{Multinomial}(|\mathcal{D}_{n}|,\mathbf{q}_{n}),$ (6)
where $|\mathcal{D}_{n}|$ denotes the local dataset size. The concentration parameter $\alpha$ directly controls the heterogeneity of the federated setting i.e. large $\alpha$ values (e.g., $\alpha\!\rightarrow\!\infty$) yield nearly IID client distributions, while small $\alpha$ values (e.g., $\alpha\!\rightarrow\!0$) produce extremely skewed, label-concentrated clients. Figure 6 visualizes this progression, showing how the distribution transitions from highly mixed to strongly partitioned as $\alpha$ decreases. We set $\alpha=0.5$ for both federated settings to maintain a balanced, overlapping non-IID distribution. For in-distribution, we choose 100 clients over a 10% participation rate, whereas two clients per domain with 100% participation rate in domain generalization setting.
[-] Base-to-New generalization. Unlike the Dirichlet distribution, we adopt a non-overlapping non-IID setting, where each client is assigned a distinct subset of classes, following the similar set up of [fedpgp]. Half of the total classes are used for training, while the remaining classes are reserved for evaluation. We configure 10 clients per dataset, with a 100% participation rate for all clients.
0.A.3 Evaluation protocol and metrics
Evaluation protocol. We evaluate the methods in the three aforementioned FL settings in a personalized FL setup [fedotp, fedpha, fedpgp], where we compute the metrics on the test set in each client and then take an average over all the clients. To be noted that client participation in training may vary according to the participation rate, but inference is performed on all clients regardless of their participation.
Metrics. In our FL experiments, we use the accuracy metric for model performance and ECE, MCE, and ACE metrics for model calibration.
[-] Accuracy. It measures the proportion of predictions for which the model’s predicted label matches the ground-truth label, providing a standard indicator of overall classification performance.
[-] Expected Calibration Error (ECE). ECE [ece] measures the average discrepancy between a model’s predicted confidence and accuracy. The prediction confidences are partitioned into $G$ static bins, and for each bin $b_{g}$, the absolute difference between the average accuracy $\text{acc}(b_{g})$ and the mean predicted confidence $\text{conf}(b_{g})$ is computed. These bin-wise errors are then weighted by the proportion of samples in each bin, yielding a global estimate of how well the model’s confidence reflects the true likelihood of correctness. Lower ECE values indicate better-calibrated predictions. The ECE metric can be defined as:
$\text{ECE}=\sum^{G}_{g=1}\frac{\mid b_{g}\mid}{|\mathcal{D}|}\mid\text{acc}(b_{g})-\text{conf}(b_{g})\mid,$ (7)
where $|\mathcal{D}|$ and $|b_{g}|$ is the total number of samples in the dataset and $g$-th bin, respectively.
[-] Maximum Calibration Error (MCE). MCE [ece] quantifies the worst-case miscalibration across all bins. Rather than averaging, MCE takes the maximum absolute gap between accuracy and confidence over the $G$ static bins. This highlights extreme local over- or under-confidence that may be masked by aggregate metrics such as ECE. A lower MCE value indicates that the model avoids severe calibration failures in any confidence region. The MCE metric can be written as:
$\text{MCE}=\max_{g\in\{1,\dots,G\}}\mid\text{acc}(b_{g})-\text{conf}(b_{g})\mid.$ (8)
[-] Adaptive Calibration Error (ACE). ACE [ace] assesses calibration by computing the discrepancy between predicted confidence and accuracy using adaptive binning. Unlike ECE which divides the predictions into $G$ static bins over fixed confidence intervals, ACE constructs $H$ adaptive bins by sorting the predicted probabilities and dividing them into bins having an equal number of samples, to provide a more realistic calculation of calibration across the entire confidence distribution. ACE evaluates calibration across all class probabilities rather than considering only the maximum predicted confidence. The ACE metric can be defined as:
$\text{ACE}=\frac{1}{C}\frac{1}{H}\sum^{C}_{c=1}\sum^{H}_{h=1}\mid\text{acc}(b_{h,c})-\text{conf}(b_{h,c})\mid,$ (9)
where $\text{acc}(b_{h,c})$ and $\text{conf}(b_{h,c})$ are respectively the average accuracy and predicted probability for class $c$ among the samples belonging to bin $h$.
0.A.4 Calibration baselines
In our experimental settings, we consider two standalone calibration methods: DCA [dca] and MDCA [mdca] as additional plug-ins in the federated methods.
[-] Difference-between-Confidence-Accuracy (DCA). DCA [dca] provides a minibatch-level calibration objective by penalizing the discrepancy between the model’s average predicted confidence score corresponding to the true class and its average accuracy for all the samples in a minibatch. For each prediction corresponding to the $i$-th sample, the method computes the correctness indicator $c_{i}$ (which equals $1$ if the prediction is correct and 0 otherwise) and the predicted confidence of the true label $s_{i}=p(\hat{y}_{i})$. By minimizing this deviation, DCA encourages the model’s overall confidence level to align with its average true accuracy, thereby improving global calibration behavior. For a minibatch of size $m$, the loss is computed as:
$\mathcal{L}_{\mathrm{dca}}=\Big\lvert\frac{1}{m}\sum_{i=1}^{m}c_{i}-\frac{1}{m}\sum_{i=1}^{m}s_{i}\Big\lvert$. (10)
Eq. 10 computes an auxiliary calibration loss by taking, over $m$ samples in a minibatch, that penalizes the mismatch between observed mean correctness and mean confidence, thereby guiding the model to better align its confidence estimates with accuracy, without a strict constraint on each example. Note that during backpropagation, the gradients are backpropagated only through the confidence term $s_{i}$.
[-] Multi-class Difference-between-Confidence-Accuracy (MDCA). MDCA [mdca] generalizes DCA to the multi-class setting by evaluating calibration across all class probabilities, rather than only the predicted class. For each class $j$, the method computes the gap between the average one-hot ground-truth indicator $c_{i}[j]$ and the average predicted probability $s_{i}[j]$ over the dataset. Averaging these discrepancies across all $C$ classes yields a calibration loss that captures class-specific trends in over- and under-confidence. MDCA thus provides a more fine-grained and stable calibration signal, which is particularly beneficial in settings with many classes or imbalanced prediction distributions. For a minibatch of size $m$,
$\mathcal{L}_{\mathrm{mdca}}$ $=\frac{1}{C}\sum_{j=1}^{C}\Big\lvert\frac{1}{m}\sum_{i=1}^{m}c_{i}[j]-\frac{1}{m}\sum_{i=1}^{m}s_{i}[j]\Big\lvert$ (11)
Eq. 11 extends the DCA concept to the multi-class setting by considering each class $j=1,\dots,C$. For each class, it computes the difference between the average ground-truth indicator $c_{i}[j]$ (which equals $1$ if sample $i$ belongs to class $j$ and 0 otherwise) and the average predicted probability $s_{i}[j]$ across all $m$ samples in a minibatch. The loss is then averaged over all $C$ classes, providing a fine-grained calibration signal that captures errors in predicted probabilities across the full class distribution.
Additional implementation details. In this paragraph, we elaborate on the implementation details, providing additional information beyond what is included in the main paper. In $\text{FL}^{2}\text{oRA}$, we only fine-tune the LoRA-adapted layers of the transformer blocks, while keeping all original weights frozen. Specifically, for both the image and text encoders, only the LoRA matrices $W_{A}$ and $W_{B}$ of the query, key, and value projections in the attention modules are trainable. All other parameters, including projection, MLP, and layer normalization weights, remain frozen to reduce memory footprint and improve training efficiency. We keep the rank of LoRA metrices to 2 and a dropout rate of 0.25 to regularize the input of the LoRA modules. Furthermore, we set the number of bins to $G=10$ for all the calibration metrics.
Appendix 0.B Additional empirical evaluation
In this section, we provide additional experimental analyses and a detailed breakdown of the results presented in the main paper.
Table 9: Comparison of federated CLIP fine-tuning methods. Results have been reported on the in-distribution setting using datasets: CIFAR-10 and CIFAR-100. Best number is in bold, and second best is underlined.
Method FL setting (w/o calib) FL setting (w/ calib: DCA [dca]) FL setting (w/ calib: MDCA [mdca])
Brier Score $\downarrow$ NLL $\downarrow$ Brier Score $\downarrow$ NLL $\downarrow$ Brier Score $\downarrow$ NLL $\downarrow$
CIFAR-10
ZS-CLIP [radford2019language] 0.17 0.37 0.17 0.37 0.17 0.37
PromptFL [promptfl] 0.12 0.23 0.12 0.24 0.12 0.23
FedPGP [fedpgp] 0.13 0.28 0.13 0.28 0.14 0.30
FedOTP [fedotp] 0.09 0.20 0.09 0.20 0.09 0.20
FedPHA [fedpha] 0.10 0.20 0.10 0.20 0.10 0.20
$\text{FL}^{2}\text{oRA}$ (Ours) 0.05 0.09 0.05 0.09 0.05 0.09
CIFAR-100
ZS-CLIP [radford2019language] 0.48 1.40 0.48 1.40 0.48 1.40
PromptFL [promptfl] 0.36 0.93 0.36 0.92 0.36 0.92
FedPGP [fedpgp] 0.42 1.12 0.38 1.00 0.38 0.98
FedOTP [fedotp] 0.38 1.00 0.37 0.97 0.38 1.00
FedPHA [fedpha] 0.35 0.91 0.34 0.90 0.34 0.91
$\text{FL}^{2}\text{oRA}$ (Ours) 0.50 1.43 0.23 0.52 0.23 0.52
0.B.1 Additional calibration metrics
Following previous work on model calibration [minderer2021revisiting], we also report performance using two additional calibration metrics, Brier score and negative log-likelihood (NLL), which complement the three other metrics reported in the main paper. First, we describe the metrics, and then we present the results.
Brier score measures the mean squared error between predicted probabilities and one-hot labels. Given $|\mathcal{D}|$ examples and $C$ classes, with $i$ indexing examples and $c$ indexing classes, let $y(i,c)\in\{0,1\}$ denote the one-hot label (1 if $c=y(i)$, else 0) and $p(i,c)\in[0,1]$ the predicted probability with $\sum_{c=1}^{C}p(i,c)=1$. It can be defined as:
$\text{Brier score}=\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\sum_{c=1}^{C}\left(p_{i,c}-y_{i,c}\right)^{2},$ (12)
Lower Brier scores indicate better calibration, as the metric penalizes assigning probability mass to incorrect classes and rewards well-calibrated, appropriately uncertain predictions.
Negative log-likelihood (NLL) evaluates the average log probability assigned to the true class and strongly penalizes overconfident mistakes. Using the same notation above, with true class index $y$, it is defined as,
$\text{NLL}=-\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\log\!\left(p_{i,y}\right),$ (13)
Lower NLL corresponds to better calibration under the model’s predictive distribution. As an unbounded, strictly proper scoring rule aligned with maximum-likelihood training, NLL is particularly sensitive to the tails of the confidence distribution, distinguishing models that differ mainly in how confidently they handle ambiguous samples.
In Table 9, we present the detailed Brier score and NLL calibration results of $\text{FL}^{2}\text{oRA}$ and the baseline methods. We observe that $\text{FL}^{2}\text{oRA}$ consistently achieve better calibration compared to the baselines. This is consistent with the findings in the main paper, where for all the calibration metrics, $\text{FL}^{2}\text{oRA}$ is better and/or competitive without needing extra regularization losses.
(a) Accuracy
(b) ECE
Figure 7: Effect of varying participation rates. We report (a) Accuracy and (b) ECE metrics on OfficeHome dataset.
0.B.2 Impact of client participation rate
In Figure 7, we showcase the effect of participation rates of clients on accuracy and ECE for CIFAR-10 and CIFAR-100 datasets. It indicates that increasing client participation generally improves accuracy for all federated methods, with gains tapering at higher rates, while ZS-CLIP has no effect of participation rates. However, our proposed $\text{FL}^{2}\text{oRA}$ consistently achieves the top accuracy across all participation levels and retains a clear margin even under sparse participation. Similarly, the calibration trends follow a similar pattern, where ECE metric is broadly stable or slightly decreases as participation rises. $\text{FL}^{2}\text{oRA}$ attains the lowest and most stable ECE, whereas alternatives such as PromptFL and FedPGP show mild, non-monotonic fluctuations.
0.B.3 Detailed results of the main tables
In Table 10, 11 and 12, we provide the detailed results of individual datasets PACS, OfficeHome and VLCS respectively in the domain-generalization setting. The average results of them are reported in the Table. 4 of the main paper. Similarly in Table 13, 14, 15, 16 and 17, we provide the detailed results of Food101, DTD, Caltech101, Flowers102, and OxfordPets datasets, corresponding to Table 5 of the main paper, for the base-to-new generalization setting.
Table 10: Comparison of federated CLIP fine-tuning methods. Performances have been reported on the domain generalization setting using the PACS dataset.
Art Painting Cartoon Photo
Method Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
ZS-CLIP [radford2019language] 98.82 3.22 2.13 3.12 93.91 2.84 1.43 2.65 99.28 3.62 1.23 3.60
FL setting (without calibration)
PromptFL [promptfl] 98.83 0.95 0.40 0.48 99.20 0.98 0.42 0.31 99.89 0.55 0.24 0.38
FedPGP [fedpgp] 99.02 0.95 0.40 0.61 99.66 0.76 0.31 0.55 99.95 0.37 0.13 0.24
FedOTP [fedotp] 98.34 4.04 1.62 3.98 99.04 3.68 1.55 3.58 99.80 2.22 1.38 2.15
FedPHA [fedpha] 99.02 0.98 0.26 0.44 99.82 0.76 0.34 0.53 99.90 0.31 0.27 0.31
$\text{FL}^{2}\text{oRA}$ (Ours) 99.41 0.57 0.20 0.36 99.93 0.69 0.28 0.42 99.85 0.49 0.18 0.23
FL setting (with calibration: DCA [dca])
PromptFL [promptfl] 98.68 1.15 0.42 0.70 99.27 1.15 0.40 0.52 99.89 0.75 0.35 0.42
FedPGP [fedpgp] 99.17 0.94 0.30 0.53 99.66 0.70 0.26 0.44 99.95 0.37 0.14 0.25
FedOTP [fedotp] 98.53 3.42 1.38 3.37 98.89 2.78 1.33 2.68 99.79 1.78 1.10 1.70
FedPHA [fedpha] 99.17 0.92 0.24 0.85 99.78 0.70 0.20 1.19 99.82 0.42 0.25 0.25
$\text{FL}^{2}\text{oRA}$ (Ours) 99.41 0.62 0.17 0.30 99.86 0.59 0.24 0.30 99.85 0.45 0.16 0.17
FL setting (with calibration: MDCA [mdca])
PromptFL [promptfl] 98.73 0.84 0.26 0.44 99.32 0.63 0.28 0.35 99.89 0.45 0.19 0.25
FedPGP [fedpgp] 98.93 1.07 0.41 0.47 99.64 1.00 0.39 0.62 99.90 0.40 0.17 0.23
FedOTP [fedotp] 98.39 4.01 1.59 3.96 98.95 3.60 1.57 3.51 99.80 2.16 1.37 2.08
FedPHA [fedpha] 99.17 1.31 0.32 0.35 99.84 1.23 0.20 0.37 99.90 0.88 0.08 0.10
$\text{FL}^{2}\text{oRA}$ (Ours) 99.41 0.55 0.17 0.36 99.93 0.56 0.25 0.38 99.90 0.43 0.18 0.26
Sketch Average
Method Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
ZS-CLIP [radford2019language] 90.99 6.41 2.38 6.58 95.75 4.02 1.80 3.99
FL setting (without calibration)
PromptFL [promptfl] 95.17 1.25 0.39 0.97 98.27 0.93 0.36 0.53
FedPGP [fedpgp] 96.89 1.47 0.50 1.10 98.88 0.89 0.34 0.63
FedOTP [fedotp] 96.18 5.18 1.77 5.07 98.34 3.78 1.58 3.70
FedPHA [fedpha] 97.08 1.32 0.46 0.72 98.96 0.84 0.33 0.50
$\text{FL}^{2}\text{oRA}$ (Ours) 97.76 0.97 0.38 0.63 99.24 0.68 0.26 0.41
FL setting (with calibration: DCA [dca])
PromptFL [promptfl] 95.30 1.51 0.49 1.46 98.29 1.14 0.42 0.78
FedPGP [fedpgp] 96.58 1.17 0.48 0.71 98.84 0.80 0.30 0.48
FedOTP [fedotp] 96.00 4.00 1.58 3.92 98.30 3.00 1.35 2.92
FedPHA [fedpha] 97.20 1.03 0.33 0.84 98.99 0.77 0.26 0.78
$\text{FL}^{2}\text{oRA}$ (Ours) 97.76 0.81 0.32 0.60 99.22 0.62 0.22 0.34
FL setting (with calibration: MDCA [mdca])
PromptFL [promptfl] 94.82 0.90 0.23 0.42 98.19 0.71 0.24 0.37
FedPGP [fedpgp] 96.33 1.43 1.67 0.93 98.70 0.98 0.66 0.56
FedOTP [fedotp] 96.08 5.28 1.79 5.12 98.30 3.76 1.58 3.67
FedPHA [fedpha] 96.88 0.88 0.35 0.50 98.95 1.07 0.24 0.33
$\text{FL}^{2}\text{oRA}$ (Ours) 97.85 1.02 0.37 0.46 99.27 0.64 0.24 0.36
Table 11: Comparison of federated CLIP fine-tuning methods. Performances have been reported on the domain generalization setting using the OfficeHome dataset.
Art Clipart Product
Method Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
ZS-CLIP [radford2019language] 83.95 10.26 1.96 8.72 65.45 5.86 1.24 6.12 89.30 3.49 1.18 3.50
FL setting (without calibration)
PromptFL [promptfl] 87.55 3.83 1.00 3.96 75.35 3.57 1.13 3.27 93.85 2.00 0.60 1.51
FedPGP [fedpgp] 89.10 5.32 1.24 3.74 78.55 4.24 1.21 3.19 95.50 2.84 0.71 1.78
FedOTP [fedotp] 81.95 23.56 5.48 23.43 77.15 17.57 3.25 17.53 92.40 21.04 3.45 21.00
FedPHA [fedpha] 88.05 3.36 0.77 2.60 81.35 3.07 0.76 2.88 96.00 1.85 0.57 1.09
$\text{FL}^{2}\text{oRA}$ (Ours) 89.39 4.73 1.09 1.86 82.08 3.76 1.08 2.76 94.38 2.28 0.70 1.24
FL setting (with calibration: DCA [dca])
PromptFL [promptfl] 86.55 3.49 0.97 3.12 76.55 4.86 1.19 3.57 94.44 2.25 0.58 1.50
FedPGP [fedpgp] 88.79 5.44 1.15 4.34 78.48 3.84 0.95 2.72 95.92 2.04 0.50 1.64
FedOTP [fedotp] 80.95 18.76 4.07 18.70 76.43 13.07 2.62 12.90 91.92 16.66 2.69 16.64
FedPHA [fedpha] 87.70 4.22 0.93 3.82 80.80 3.64 0.87 2.96 95.93 1.94 0.52 1.99
$\text{FL}^{2}\text{oRA}$ (Ours) 88.90 2.59 0.82 3.32 82.35 2.68 0.84 2.28 93.82 2.08 0.48 1.98
FL setting (with calibration: MDCA [mdca])
PromptFL [promptfl] 86.55 4.47 1.19 2.96 75.34 3.81 0.81 3.02 94.13 1.88 0.63 1.35
FedPGP [fedpgp] 88.18 3.57 0.91 3.46 78.65 4.89 1.76 3.61 95.93 3.15 0.76 2.32
FedOTP [fedotp] 82.28 23.86 6.32 23.66 77.50 18.06 3.75 17.97 92.19 20.59 3.16 20.55
FedPHA [fedpha] 89.27 2.81 0.78 2.09 80.62 3.63 1.06 4.16 95.73 1.26 0.44 1.26
$\text{FL}^{2}\text{oRA}$ (Ours) 89.39 4.91 1.17 2.23 81.83 3.40 1.07 2.60 94.38 2.30 0.74 1.25
Real World Average
Method Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
ZS-CLIP [radford2019language] 90.25 5.15 1.47 4.96 82.24 6.19 1.46 5.82
FL setting (without calibration)
PromptFL [promptfl] 93.00 1.81 0.62 1.71 87.44 2.80 0.84 2.61
FedPGP [fedpgp] 93.95 3.51 0.84 2.17 89.28 3.98 1.00 2.72
FedOTP [fedotp] 88.00 16.38 2.65 16.38 84.88 19.64 3.71 19.59
FedPHA [fedpha] 93.50 2.32 0.80 1.64 89.73 2.65 0.73 2.05
$\text{FL}^{2}\text{oRA}$ (Ours) 93.14 1.64 0.47 1.58 89.75 3.10 0.83 1.86
FL setting (with calibration: DCA [dca])
PromptFL [promptfl] 93.76 2.24 0.69 2.26 87.83 3.21 0.86 2.61
FedPGP [fedpgp] 93.83 2.47 0.90 2.06 89.25 3.45 0.87 2.69
FedOTP [fedotp] 87.15 12.89 2.26 12.86 84.11 15.35 2.91 15.28
FedPHA [fedpha] 92.82 2.95 0.91 2.66 89.31 3.19 0.81 2.86
$\text{FL}^{2}\text{oRA}$ (Ours) 92.66 1.99 0.49 1.33 89.43 2.34 0.66 2.23
FL setting (with calibration: MDCA [mdca])
PromptFL [promptfl] 93.97 2.56 0.63 2.63 87.50 3.18 0.82 2.49
FedPGP [fedpgp] 94.40 3.23 0.77 3.12 89.29 3.71 1.05 3.13
FedOTP [fedotp] 87.64 16.21 2.92 16.21 84.90 19.68 4.04 19.60
FedPHA [fedpha] 93.93 2.40 0.69 1.79 89.89 2.53 0.74 2.33
$\text{FL}^{2}\text{oRA}$ (Ours) 93.03 1.46 0.47 1.47 89.66 3.02 0.86 1.89
Table 12: Comparison of federated CLIP fine-tuning methods. Performances have been reported on the domain generalization setting using the VLCS dataset.
Caltech Labelme Pascal-VOC
Method Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
ZS-CLIP [radford2019language] 78.44 10.36 4.56 10.53 84.65 6.40 2.41 6.64 87.94 7.51 1.75 8.24
FL setting (without calibration)
PromptFL [promptfl] 100.00 2.22 1.83 2.15 78.24 10.11 4.84 11.67 89.27 2.70 1.19 2.58
FedPGP [fedpgp] 100.00 2.30 1.23 2.25 81.37 4.87 2.09 5.68 90.77 2.06 1.07 2.30
FedOTP [fedotp] 100.00 2.80 1.82 2.80 81.96 6.20 2.13 6.08 90.58 6.21 1.80 5.59
FedPHA [fedpha] 100.00 2.51 1.18 2.39 81.98 4.85 1.92 5.40 90.67 2.11 0.88 1.82
$\text{FL}^{2}\text{oRA}$ (Ours) 100.00 2.04 1.20 2.01 81.57 5.79 1.84 5.25 91.06 2.59 1.09 2.27
FL setting (with calibration: DCA [dca])
PromptFL [promptfl] 100.00 2.28 1.78 2.23 78.34 8.90 4.69 10.27 90.61 2.77 0.82 2.59
FedPGP [fedpgp] 100.00 0.22 0.13 0.13 86.20 6.32 1.96 6.55 95.85 1.72 0.69 2.34
FedOTP [fedotp] 100.00 2.14 1.45 2.14 82.65 5.00 1.40 4.75 94.50 5.24 1.53 4.50
FedPHA [fedpha] 100.00 0.32 0.11 0.14 86.53 4.54 1.34 5.05 94.82 1.62 0.52 1.55
$\text{FL}^{2}\text{oRA}$ (Ours) 100.00 2.18 1.10 2.16 83.16 5.31 1.65 4.88 90.67 3.20 1.38 2.67
FL setting (with calibration: MDCA [mdca])
PromptFL [promptfl] 100.00 2.00 1.69 1.99 79.68 8.37 4.52 9.55 90.91 2.46 0.75 3.30
FedPGP [fedpgp] 100.00 0.27 0.21 0.21 87.91 3.26 1.31 4.62 95.59 1.63 0.69 1.36
FedOTP [fedotp] 100.00 2.85 1.72 2.85 83.89 5.87 1.58 6.32 94.37 5.89 1.67 5.45
FedPHA [fedpha] 100.00 1.98 1.77 1.79 86.64 4.05 2.51 3.92 94.72 2.66 1.61 2.66
$\text{FL}^{2}\text{oRA}$ (Ours) 99.82 1.97 1.13 1.96 82.55 5.39 2.01 4.24 91.06 2.44 0.75 2.34
Sun Average
Method Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
ZS-CLIP [radford2019language] 80.78 3.95 1.45 4.22 82.95 7.06 2.54 7.41
FL setting (without calibration)
PromptFL [promptfl] 84.63 4.53 1.89 4.45 88.03 4.89 2.44 5.21
FedPGP [fedpgp] 87.81 3.32 0.87 3.25 89.99 3.14 1.32 3.37
FedOTP [fedotp] 87.38 3.19 1.46 2.85 89.98 4.60 1.80 4.33
FedPHA [fedpha] 87.06 4.04 1.49 3.67 89.93 3.38 1.37 3.32
$\text{FL}^{2}\text{oRA}$ (Ours) 88.77 2.88 1.01 3.31 90.35 3.33 1.29 3.21
FL setting (with calibration: DCA [dca])
PromptFL [promptfl] 82.49 5.51 2.23 5.84 87.86 4.87 2.38 5.23
FedPGP [fedpgp] 92.79 2.10 0.71 2.96 93.71 2.59 0.87 3.00
FedOTP [fedotp] 91.62 2.67 1.34 3.10 92.19 3.76 1.43 3.62
FedPHA [fedpha] 92.59 3.69 1.44 3.00 93.49 2.54 0.85 2.44
$\text{FL}^{2}\text{oRA}$ (Ours) 89.17 2.61 1.01 3.41 90.75 3.33 1.29 3.28
FL setting (with calibration: MDCA [mdca])
PromptFL [promptfl] 82.49 6.59 3.37 6.96 88.27 4.85 2.58 5.45
FedPGP [fedpgp] 93.08 3.54 1.73 1.79 94.14 2.17 0.98 2.00
FedOTP [fedotp] 91.45 2.72 1.31 3.22 92.43 4.33 1.57 4.46
FedPHA [fedpha] 91.99 3.69 1.48 3.01 93.34 3.10 1.84 2.85
$\text{FL}^{2}\text{oRA}$ (Ours) 87.71 2.44 0.90 2.13 90.28 3.06 1.20 2.67
Table 13: Comparison of federated CLIP fine-tuning methods. Performances have been reported on the base-to-new generalization setting using the Food101 dataset.
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
Method Base New HM Base New HM Base New HM Base New HM
ZS-CLIP [radford2019language] 90.29 91.70 90.99 2.38 2.90 2.61 0.76 0.82 0.79 2.38 2.90 2.61
FL setting (without calibration)
PromptFL [promptfl] 89.78 90.79 90.28 0.48 0.70 0.57 0.13 0.17 0.15 0.39 0.61 0.47
FedPGP [fedpgp] 87.54 87.94 87.74 1.19 1.08 1.13 0.38 0.35 0.37 1.16 1.02 1.09
FedOTP [fedotp] 15.18 40.18 22.03 39.93 17.49 24.32 8.05 4.02 5.36 39.93 17.48 24.31
FedPHA [fedpha] 58.11 70.69 63.79 19.21 7.56 10.85 7.37 2.78 4.04 18.19 7.31 10.43
$\text{FL}^{2}\text{oRA}$ (Ours) 89.34 89.83 89.58 1.95 1.12 1.42 0.62 0.29 0.39 1.81 1.03 1.31
FL setting (with calibration: DCA [dca])
PromptFL [promptfl] 89.87 91.11 90.49 0.57 0.89 0.70 0.10 0.30 0.15 0.58 0.83 0.68
FedPGP [fedpgp] 85.46 86.16 85.81 2.72 0.99 1.45 0.99 0.33 0.50 2.68 0.92 1.37
FedOTP [fedotp] 14.08 38.27 20.59 44.94 20.15 27.83 8.26 5.34 6.49 44.93 20.13 27.80
FedPHA [fedpha] 54.03 69.22 60.69 25.12 9.76 14.06 9.79 3.13 4.74 23.47 9.40 13.42
$\text{FL}^{2}\text{oRA}$ (Ours) 89.21 89.82 89.51 2.46 1.46 1.83 0.81 0.38 0.52 2.25 1.37 1.70
FL setting (with calibration: MDCA [mdca])
PromptFL [promptfl] 89.81 90.60 90.20 0.39 0.81 0.53 0.11 0.27 0.16 0.45 0.77 0.57
FedPGP [fedpgp] 87.54 87.94 87.74 1.19 1.08 1.13 0.38 0.35 0.36 1.16 1.02 1.08
FedOTP [fedotp] 18.91 40.78 25.84 35.79 13.38 19.48 7.18 2.84 4.07 35.79 13.38 19.48
FedPHA [fedpha] 50.77 60.46 55.19 26.35 17.71 21.18 10.07 6.04 7.55 24.84 17.14 20.28
$\text{FL}^{2}\text{oRA}$ (Ours) 89.35 89.81 89.58 2.01 1.17 1.48 0.64 0.30 0.41 1.86 1.10 1.38
Table 14: Comparison of federated CLIP fine-tuning methods. Performances have been reported on the base-to-new generalization setting using the DTD dataset.
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
Method Base New HM Base New HM Base New HM Base New HM
ZS-CLIP [radford2019language] 53.70 54.23 53.96 5.90 8.47 6.96 1.49 2.97 1.98 6.85 9.88 8.09
FL setting (without calibration)
PromptFL [promptfl] 76.85 44.93 56.71 6.72 24.65 10.56 2.32 7.38 3.54 6.35 24.65 10.10
FedPGP [fedpgp] 67.88 50.87 58.16 13.98 19.19 16.18 5.88 4.27 4.95 13.88 19.12 16.08
FedOTP [fedotp] 27.65 42.17 33.40 41.26 13.68 20.55 11.00 2.95 4.66 41.04 13.71 20.55
FedPHA [fedpha] 27.87 38.29 32.26 50.52 20.53 29.20 25.81 4.50 7.67 48.96 20.50 28.90
$\text{FL}^{2}\text{oRA}$ (Ours) 82.63 55.59 66.47 8.45 26.27 12.79 3.12 10.45 4.81 7.09 23.35 10.88
FL setting (with calibration: DCA [dca])
PromptFL [promptfl] 78.59 51.21 62.01 6.78 20.79 10.23 1.93 5.32 2.83 6.41 20.79 9.80
FedPGP [fedpgp] 69.84 53.85 60.81 10.94 12.93 11.85 3.65 2.62 3.05 10.77 12.85 11.72
FedOTP [fedotp] 22.66 42.13 29.47 50.82 16.65 25.09 16.58 4.04 6.50 50.52 16.70 25.11
FedPHA [fedpha] 25.81 39.55 31.24 53.81 20.70 29.90 27.74 5.39 9.03 51.43 20.44 29.26
$\text{FL}^{2}\text{oRA}$ (Ours) 82.85 55.98 66.81 9.91 26.80 14.47 4.08 11.01 5.95 8.26 23.32 12.20
FL setting (with calibration: MDCA [mdca])
PromptFL [promptfl] 76.62 53.86 63.26 7.17 17.06 10.10 2.37 4.10 3.00 6.71 16.75 9.58
FedPGP [fedpgp] 74.72 48.54 58.85 7.88 19.95 11.30 3.34 3.91 3.60 7.63 19.85 11.02
FedOTP [fedotp] 29.10 44.70 35.25 39.73 10.15 16.17 11.06 2.19 3.66 39.45 10.13 16.12
FedPHA [fedpha] 26.47 37.40 31.00 52.05 23.25 32.14 25.88 5.45 9.01 50.00 23.00 31.51
$\text{FL}^{2}\text{oRA}$ (Ours) 82.47 56.11 66.78 9.05 26.77 13.53 3.74 10.69 5.54 7.69 23.97 11.64
Table 15: Comparison of federated CLIP fine-tuning methods. Performances have been reported on the base-to-new generalization setting using the Caltech101 dataset.
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
Method Base New HM Base New HM Base New HM Base New HM
ZS-CLIP [radford2019language] 95.93 93.34 94.62 7.41 3.34 4.60 1.50 2.55 1.89 7.41 2.68 3.94
FL setting (without calibration)
PromptFL [promptfl] 97.55 92.36 94.88 0.77 4.41 1.31 0.21 3.77 0.40 0.36 4.89 0.67
FedPGP [fedpgp] 96.36 92.43 94.36 1.39 4.08 2.07 0.62 3.01 1.03 0.87 3.64 1.40
FedOTP [fedotp] 38.97 67.98 49.54 18.51 11.40 14.11 3.68 2.61 3.05 18.68 11.31 14.09
FedPHA [fedpha] 82.15 87.03 84.52 5.42 5.55 5.48 1.95 2.75 2.28 5.16 4.75 4.95
$\text{FL}^{2}\text{oRA}$ (Ours) 98.34 93.41 95.81 1.34 3.33 1.91 0.44 1.17 0.64 0.84 2.51 1.26
FL setting (with calibration: DCA [dca])
PromptFL [promptfl] 97.09 92.03 94.49 1.56 4.87 2.36 0.89 4.24 1.47 1.08 4.29 1.73
FedPGP [fedpgp] 96.95 92.98 94.93 1.01 3.89 1.60 0.41 2.75 0.71 0.82 3.26 1.31
FedOTP [fedotp] 42.63 66.62 51.99 19.08 7.75 11.02 3.87 1.95 2.59 19.13 7.84 11.12
FedPHA [fedpha] 78.78 87.05 82.71 8.54 6.13 7.14 4.03 3.06 3.48 8.05 5.04 6.20
$\text{FL}^{2}\text{oRA}$ (Ours) 98.30 93.33 95.75 1.20 3.44 1.78 0.39 1.05 0.57 0.60 2.67 0.98
FL setting (with calibration: MDCA [mdca])
PromptFL [promptfl] 97.03 92.58 94.75 1.20 4.40 1.89 0.37 3.35 0.67 0.55 3.58 0.95
FedPGP [fedpgp] 96.45 92.45 94.40 1.30 2.63 1.74 0.65 1.09 0.81 1.00 2.40 1.41
FedOTP [fedotp] 31.80 68.20 43.38 26.03 12.32 16.72 5.35 2.95 3.80 25.99 12.23 16.63
FedPHA [fedpha] 78.68 84.12 81.31 7.86 6.60 7.18 2.86 2.77 2.81 7.46 5.39 6.26
$\text{FL}^{2}\text{oRA}$ (Ours) 98.36 93.43 95.83 1.30 3.35 1.87 0.43 1.23 0.64 0.83 2.49 1.24
Table 16: Comparison of federated CLIP fine-tuning methods. Performances have been reported on the base-to-new generalization setting using the Flowers102 dataset.
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
Method Base New HM Base New HM Base New HM Base New HM
ZS-CLIP [radford2019language] 73.22 77.94 75.51 4.83 4.45 4.63 1.38 0.94 1.12 5.20 4.89 5.04
FL setting (without calibration)
PromptFL [promptfl] 95.82 64.26 76.93 3.05 15.74 5.11 1.04 3.89 1.64 2.91 15.74 4.91
FedPGP [fedpgp] 73.79 65.21 69.24 7.53 10.88 8.90 3.24 2.66 2.92 7.41 10.86 8.81
FedOTP [fedotp] 15.15 32.85 20.73 46.74 13.96 21.50 9.81 3.28 4.91 46.72 14.09 21.65
FedPHA [fedpha] 29.43 49.37 36.88 48.27 16.31 24.38 21.26 3.57 6.11 45.80 16.28 24.02
$\text{FL}^{2}\text{oRA}$ (Ours) 98.01 72.27 83.19 1.44 15.23 2.63 0.46 5.72 0.85 1.02 14.47 1.90
FL setting (with calibration: DCA [dca])
PromptFL [promptfl] 94.68 66.74 78.29 2.69 11.13 4.33 0.97 2.16 1.33 2.32 11.14 3.84
FedPGP [fedpgp] 70.77 64.21 67.33 11.15 14.53 12.62 4.90 3.79 4.27 11.04 14.52 12.54
FedOTP [fedotp] 14.01 28.67 18.82 53.06 22.78 31.88 13.38 5.57 7.86 53.06 22.93 32.02
FedPHA [fedpha] 29.05 47.58 36.08 50.91 20.96 29.69 22.98 4.47 7.49 47.93 20.96 29.17
$\text{FL}^{2}\text{oRA}$ (Ours) 97.83 70.65 82.05 1.07 16.88 2.01 0.34 6.55 0.65 0.60 15.82 1.16
FL setting (with calibration: MDCA [mdca])
PromptFL [promptfl] 95.16 67.30 78.84 3.32 11.70 5.18 1.07 2.77 1.54 3.15 11.71 4.96
FedPGP [fedpgp] 74.97 66.35 70.40 7.59 9.26 8.34 3.02 2.24 2.57 7.63 9.32 8.39
FedOTP [fedotp] 13.92 24.50 17.75 51.17 26.85 35.22 11.34 6.65 8.39 51.16 26.75 35.13
FedPHA [fedpha] 29.01 45.74 35.50 49.52 21.80 30.27 22.22 5.88 9.30 46.80 21.70 29.65
$\text{FL}^{2}\text{oRA}$ (Ours) 98.02 72.49 83.35 1.37 15.13 2.51 0.43 5.70 0.80 0.98 14.36 1.83
Table 17: Comparison of federated CLIP fine-tuning methods. Performances have been reported on the base-to-new generalization setting using the OxfordPets dataset.
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
Method Base New HM Base New HM Base New HM Base New HM
ZS-CLIP [radford2019language] 92.88 97.60 95.18 4.39 3.36 3.81 1.36 1.35 1.36 4.27 3.37 3.77
FL setting (without calibration)
PromptFL [promptfl] 95.48 96.76 96.12 2.85 3.13 2.98 0.79 0.91 0.85 2.78 3.13 2.95
FedPGP [fedpgp] 94.86 97.12 95.98 2.08 3.37 2.57 0.86 1.42 1.07 2.00 3.25 2.48
FedOTP [fedotp] 13.86 52.07 21.89 64.69 18.64 28.94 26.50 4.13 7.15 64.55 18.63 28.91
FedPHA [fedpha] 44.75 80.04 57.40 36.96 5.96 10.26 20.54 1.69 3.12 33.76 5.96 10.13
$\text{FL}^{2}\text{oRA}$ (Ours) 95.16 97.29 96.22 0.60 1.24 0.81 0.33 0.41 0.37 0.66 0.69 0.67
FL setting (with calibration: DCA [dca])
PromptFL [promptfl] 95.37 96.59 95.98 2.25 2.19 2.22 0.65 0.75 0.70 2.24 1.90 2.06
FedPGP [fedpgp] 94.05 97.21 95.60 1.11 2.78 1.59 0.38 1.06 0.56 1.08 2.64 1.53
FedOTP [fedotp] 14.33 50.29 22.31 63.71 26.13 37.06 27.50 8.75 13.28 63.60 26.19 37.10
FedPHA [fedpha] 51.62 77.82 62.07 29.78 5.22 8.88 17.39 1.60 2.93 27.62 5.27 8.85
$\text{FL}^{2}\text{oRA}$ (Ours) 95.35 97.22 96.28 0.98 0.93 0.95 0.42 0.30 0.35 0.69 0.48 0.57
FL setting (with calibration: MDCA [mdca])
PromptFL [promptfl] 95.37 96.76 96.06 2.59 2.64 2.61 0.72 0.87 0.79 2.55 2.64 2.59
FedPGP [fedpgp] 95.02 97.29 96.14 1.85 2.60 2.16 0.80 1.02 0.90 1.77 2.56 2.09
FedOTP [fedotp] 13.08 46.77 20.45 64.72 28.09 39.18 25.73 7.28 11.35 64.66 28.19 39.26
FedPHA [fedpha] 46.56 79.30 58.67 33.12 7.30 11.96 18.36 2.56 4.49 30.86 7.14 11.60
$\text{FL}^{2}\text{oRA}$ (Ours) 95.10 97.28 96.17 1.03 1.18 1.10 0.38 0.39 0.39 0.72 0.63 0.67